# David Roodman Open Philanthropy San Francisco, CA

### v2

@@ 3.1 Estimation with high-dimensional fixed effects
The introduction of reghdfe in 2014 (Correia, 2016) was an important event in applied econometrics. Building on Guimarães and Portugal (2010) and Gaure (2011), the Mata-based program proved the computational tractability of linear models with many sets of fixed effects. Since then, other authors have developed alternative algorithms for absorbing fixed effects, several of which have been incorporated into reghdfe. The ideas have been extended to the Poisson model through ppmlhdfe for Stata (Correia, Guimarães, and Zylkin, 2019), as well as to other generalized linear models, in alpaca and fixest for R (Stammann 2018; Bergé 2018). Julia packages for HDFE modeling are also available: Matthieu Gomez’s FixedEffectModels.jl and Johannes Boehm’s GLFixedEffectModels.jl. With regard to linear models, according to benchmarking results on the GitHub pages of fixest and FixedEffectModels.jl, those two packages are 1–2 orders of magnitude faster than regdhfe. [^7] See github.com/lrberge/fixest#benchmarking and github.com/FixedEffects/FixedEffectModels.jl#benchmarks.
These developments have given R and Julia users faster and more general implementations of HDFE modeling than were available in Stata. That is what motivated the development of the julia and reghdfejl packages. The idea was to not reinvent wheels, but to fashion a familiar-looking chassis onto existing wheels.
@@
In spirit, reghdfejl is a slot-in replacement for reghdfe, with the same syntax, output, and e() return values. In practice, the two differ outwardly as well as inwardly. As reghdfjle depends on the feature set of FixedEffectModels.jl, it offers only one optimization technique, LSMR (Fong and Saunders, 2011)—though this has never been a problem in my experience. reghdfejl does not support “group fixed effects.” And it does not adjust the reported degrees of freedom for any collinearity of the fixed effects with each other or with the error clustering. So it often reports slightly larger standard errors. While reghdfejl can perform instrumental variables (IV) estimation, unlike ivregdfe it does not by default work as a wrapper for ivreg2 (Baum, Schaffer, and Stillman, 2007). As a result, it can only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic. reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and pass the results to ivreg2. But in this mode, reghdfejl is no faster than ivreghdfe. For on hard problems, ivreg2 dominates the runtime.
reghdfe and reghdfejl share the following options: absorb(), vce(), residuals(), tolerance(), iterate(#), nosample, keepsingletons, compact, level(#), and standard Stata regression display options. In particular, reghdfejl’s absorb() can save some or all of the estimated fixed effects. For instance, a(firm year_fe=year) will absorb firm and year fixed effects, and store estimates of the latter in the variable year_fe. For both reghdfe and reghdfejl, the tolerance() option sets a precision threshold for the fitting algorithms to declare convergence; however, the semantics are not necessarily the same, so that, say, tol(1e-6) could elicit slightly different results from the two.
reghdfejl accepts three distinctive (documented) options:
@@
threads(#) lowers the limit on the number of threads that can be used. It cannot increase the thread count above the limit set when Julia is started (see section 4.2). This option is rarely used.
verbose directs reghdfejl to show more of its work—to display the Julia copy of the data set, the formula for the regression model, and the regression command. The data set and formula are left behind for the user to work with through jl, not erased as they otherwise would be.
@@
specifies that 1000 bootstrap replications be performed, split across four copies of Julia running in parallel, with the bootstrap data drawn groupwise from clusters defined by year. One could attain the same results with bootstrap:reghdfe or bs:reghdfe. But reghdfejl implements vce(bs) in a radically faster way. After launching the extra instances of Julia to split the work, it copies the primary data just once to each of them, and effects the sampling in Julia by updating a weight variable (cluster 1’s weights are doubled if drawn two times, etc.).
To compare reghdfe and reghdfejl on speed, I create a data set with 10 million rows and two sets of fixed effects. I run the same regression with both commands. I explore the effect of switching from one to multiple processors in Stata/MP, and of adding reghdfejl’s gpu option:
@@
    reghdfe   y x1 x2, a(id1 id2) cluster(id1 id2) dof(none)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2) gpu
@@
    reghdfe   y x1 x2, a(id1 id2) cluster(id1 id2) dof(none)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1 id2) gpu
For comparability, the reghdfe lines include dof(none) to disable the degrees-of-freedom correction, since reghdfejl lacks that feature. The first triplet runs on one CPU core, as under a non-MP flavor of Stata. But that setting on restricts the ado and Mata code: the Julia code still runs on multiple CPU threads. On a laptop with an Intel i9-13900H CPU and an NVIDIA RTX 4070 GPU, reghdfe takes 32 seconds. reghdfejl takes 3.8 seconds without the gpu option and 2.8 with. Moving to 6 processors in Stata and Mata—since the Intel chip has 6 “performance” cores—lowers the times to 24, 3.3, and 2.7 seconds.
@@
Using Stata’s bs prefix command with reghdfe takes 64 seconds. Switching to the parallel package of George Vega and Brian Quistorff in order to split the work among 6 copies of Stata [^8] See github.com/gvegayon/parallel. cuts the time to 23 seconds. reghdfejl takes just 2.5 seconds. But, as usual, the Julia function is slower the first time it is invoked in a Stata session: this one spins up 6 worker copies of Julia and triggers some just-in-time compilation in each. The workers remain active after completing their first task, so subsequent bootstrapping is fast as long as the number of workers is kept at 6.
The reghdfejl package includes a partialhdfejl command for absorbing fixed effects out of other variables. partialhdfejl command lines must contain either a generate() option to name the variables that will hold the results, or a prefix() option for making the new variable names out of the old. If any of the new variables already exists, partialhdfejl will error, unless the replace option is also included. The following two runs produce the same point estimates:
@@ 3.2 Wild bootstrapping
Roodman et al. (2019) introduces the Stata package boottest for inference via the wild bootstrap. Among the methods boottest implements is the “wild restricted efficient” bootstrap (WRE), which is the Davidson and MacKinnon (2010) adaptation for instrumental variables (IV) estimation. Since the publication of Roodman et al. (2019), boottest has been optimized in ways that especially benefit the WRE. Using the Frisch-Waugh-Lowell theorem, exogenous controls are now partialled out of the other variables using the same algebraic tricks that accelerate boottest’s wild bootstrap for ordinary least squares. The partialling-out accelerates the application of the user’s IV estimator despite its irreducible nonlinearity (Roodman et al., 2019, section 6.1). Also incorporated (more fully), is MacKinnon (2023)’s point that the data only enter wild-bootstrapped inferences through indexed sets of matrix products such as $\{{\bf X}^{\prime}_{g}{\bf Y}_{g}:g=1,...,G\}$ where ${\bf X}_{g}$ and ${\bf Y}_{g}$ hold the rows of data matrices ${\bf X}$ and ${\bf Y}$ for cluster $g$. Reducing the data to these products early in the calculation complicates the algebra the code must implement, but speeds execution when clusters are few.
Finally, the back-end functionality in boottest has been transplanted to the Julia package WildBootTests.jl. Adding the julia option to a boottest command line switches to the new back end.
Section 8.4 of Roodman et al. (2019) approximately replicates two 2SLS regressions in the Levitt (1996) study of the short-term impact of decarceration on crime. [^9] Data and preparation code are available at davidroodman.com/david/Levitt.zip. Here is a version of that example that substitutes reghdfejl for ivregress and sets the stage for adding the julia option to the boottest command line:
@@
On a single CPU core, version 2.3.5 of boottest, the one published with Roodman et al. (2019), needs 210 seconds to perform the two boottest calls. The latest version finishes in 7.4 seconds. Adding the julia option to the boottest line brings the time to 4.1 seconds.
The R package fwildclusterboot (Fischer and Roodman, 2021) also calls WildBootTests.jl to perform the WRE. This may be the first instance of a Julia statistical package serving as a back end on multiple platforms.
@@ 3.3 Plotting
Many data visualization packages are available for Julia. Surveying them is beyond the scope of this article. To indicate their potential for Stata users, I generate two graphs with Makie.jl.
Makie works with several plotting back ends, which render output in formats that are optimized for different applications, such as interactive graphs in web pages and publication-quality image files. The Cairo back end is best for the latter. It is bundled with Makie in CairoMakie.jl:
  • jl AddPkg CairoMakie  // one-time installation
    jl: using CairoMakie   // use now
In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.101010blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command. margins is then called to compute the probability of high blood pressure according to the fitted model, across a grid of ages and weights. The probabilities are copied to Python and depicted in a three-dimensional surface plot. A Julia adaptation of that example is:
@@
This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window, That is, I drop the “;” and “///” from the ends of lines. The first Julia line imports the data as a matrix and then converts it to a DataFrame with named columns. Then the surface plot is built and stored in the Julia variable f. The call to surface() refers to columns of the DataFrame with the convenient dot syntax, as in df.age; here, column names are not prefixed with a colon. The next command displays the figure in a new window, which might pop up behind the Stata window. That line takes advantage of Julia’s piping syntax for function calls. Last, the image is saved to disk. It appears in Figure 1.
@@
dropmissing!(df, :bmi) is similar to drop if bmi==. in Stata. However, it changes the type of df.bmi from Float32? to Float32, meaning that the column no longer accepts missing values. That is apparently necessary to make Makie work. The “!” in “dropmissing!” has no special meaning in the syntax of Julia: it is just a character in the function’s name. By convention, functions that modify their arguments have names ending in “!”.
The for...enumerate() loop iterates over the racial groups. r takes race names as string values while i is a counter starting from 1. In the Stata data set, the race column is coded with 1, 2, and 3. Those codes are displayed through a value label as “White”, “Black”, and “Other.” Since jl save stores race as a CategoricalVector, the Julia column’s contents are outwardly strings, not numbers. df.bmi[df.race .== r] uses Boolean indexing to extract the elements of the bmi column for which race takes the string value in r.
Each iteration plots a density in slate gray with 40% opacity. The spacing offset of $1/20$ is interpreted in the vertical axis’s units, BMI density, and was chosen by trial and error to cause overlapping. See Figure 2.
@@ 4.1 The JuliaMono font
Julia output routinely contains certain Unicode characters, such as vertical and diagonal ellipses, that Stata fonts do not cover. (However, this is not necessarily the case for Stata packages that call Julia back ends.) The JuliaMono font is a free, monospaced font that includes those characters, and looks reasonably good in Stata’s results window. The font can be installed from github.com/cormullion/juliamono/releases. Type help fonts in Stata to learn how to change the Results window font.
@@ 4.2 Multithreading
Julia includes several facilities for multitasking. One is distributed processing, which involves launching multiple copies of Julia that can pass messages and data to each other. reghdfejl’s vce(bs) option uses that feature. A lighter form is multithreading, in which multiple execution lines run within one Julia process, largely sharing the same data space. That is the sort @tturbo exploits. The maximum number of threads Julia allows is set when it launches, and it can default to 1. A good guess for the optimum is the number of “performance” or “big” cores on the CPU. You can determine the current limit with the Julia command Threads.nthreads().
How best to set Julia’s thread limit depends on the operating system. In Windows, you can open the Environment Variables control panel and assign JULIA_NUM_THREADS. In Linux, you can add an entry such as export JULIA_NUM_THREADS = 6 to $\sim$/.bashrc” or the equivalent. In macOS, environment variables do not affect graphical applications such as Stata when launched from the desktop. One can instead set the limit in each Stata session through the threads() option of the jl start command. Example: jl start, threads(8). This must be the first jl command issued since starting Stata. It must also precede invocation of Julia-dependent programs such as reghdfejl. [^11] In macOS, one can in fact assign JULIA_NUM_THREADS in $\sim$/.zshenv”. This will take effect when one launches Stata from the terminal. For example open -a StataMP will launch Stata/MP.
4.3 Limitations on accessing Stata objects
The julia package lets you transfer information between Stata and Julia with commands issued in either language; section 5 will lay out more details. However initiated, the job of copying falls to the C++ plugin. To read and write Stata macros, scalars, matrices, and variables, the plugin must in turn call the official Stata plugin interface. [^12] See stata.com/plugins. The SPI imposes certain constraints. For one, it makes public no functions that can create Stata variables or matrices. As a result, you can edit existing Stata variables and matrices from Julia, but not create new ones. The Stata-side jl subcommands can and do evade this constraint because they start from the Stata environment. For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy from a Julia matrix.
A subtler limitation pertains to local macros. Explaining it requires a deeper dive. The “jl.ado” file defines the jl command. In order to call the plugin “jl.plugin,” it first declares the plugin’s existence with this line:

### v3

@@ 3.1 Estimation with high-dimensional fixed effects
The introduction of reghdfe in 2014 (Correia, 2016) was an important event in applied econometrics. Building on Guimarães and Portugal (2010) and Gaure (2011), the Mata-based program proved the computational tractability of linear models with many sets of fixed effects. Since then, other authors have developed alternative algorithms for absorbing fixed effects, several of which have been incorporated into reghdfe. The ideas have been extended to the Poisson model through ppmlhdfe for Stata (Correia, Guimarães, and Zylkin, 2019), as well as to other generalized linear models, in alpaca and fixest for R (Stammann 2018; Bergé 2018). Julia packages for HDFE modeling are also available: Matthieu Gomez’s FixedEffectModels.jl and Johannes Boehm’s GLFixedEffectModels.jl. According to benchmarking results on the GitHub pages of fixest and FixedEffectModels.jl, those two packages are 1–2 orders of magnitude faster than regdhfe. [^7] See github.com/lrberge/fixest#benchmarking and github.com/FixedEffects/FixedEffectModels.jl#benchmarks.
These developments have given R and Julia users faster and more general implementations of HDFE modeling than have been available in Stata. That is what motivated the development of the julia and reghdfejl packages. The idea was to not reinvent wheels, but to fashion a familiar-looking Stata chassis onto existing Julia wheels.
@@
In spirit, reghdfejl is a slot-in replacement for reghdfe, with the same syntax, output, and e() return values. In practice, the two differ outwardly as well as inwardly. As reghdfjl depends on the feature set of FixedEffectModels.jl, it offers only one optimization technique, LSMR (Fong and Saunders, 2011)—though this has never been a problem in my experience. reghdfejl does not support “group fixed effects.” And it does not adjust the reported degrees of freedom for any collinearity of the fixed effects with each other or with the error clustering. So it sometimes reports slightly larger standard errors. While reghdfejl can perform instrumental variables (IV) estimation, unlike ivregdfe it does not by default work as a wrapper for ivreg2 (Baum, Schaffer, and Stillman, 2007). As a result, it can only perform two-stage least squares (2SLS); and for weak identification diagnosis, it only reports the Kleibergen-Paap first-stage F statistic. reghdfejl does accept an ivreg2 option, which directs the program to call FixedEffectModels.jl to absorb the fixed effects and then transfer control to ivreg2. But in this mode, reghdfejl is no faster than ivreghdfe. For on hard problems, ivreg2 accounts for most of the run time.
reghdfe and reghdfejl share the following options: absorb(), vce(), residuals(), tolerance(), iterate(#), nosample, keepsingletons, compact, level(#), and standard Stata regression display options. In particular, reghdfejl’s absorb() can save some or all of the estimated fixed effects. For instance, a(firm year_fe=year) will absorb firm and year fixed effects, and store estimates of the latter in the variable year_fe. For both reghdfe and reghdfejl, the tolerance() option sets a precision threshold for the fitting algorithms to declare convergence; however, the semantics are not necessarily the same, so that, say, tol(1e-6) could elicit slightly different results from the two programs.
reghdfejl accepts three distinctive options:
@@
threads(#) caps the number of parallel threads that can be used. It cannot increase the thread count above the limit set when Julia is started (see section 4.2). This option is rarely used.
verbose directs reghdfejl to show more of its work—to display the Julia copy of the data set, the formula for the regression model, and the Julia regression command. The data set and formula are left behind for the user to work with  not erased as they otherwise would be.
@@
specifies that 1000 bootstrap replications be performed, split across four copies of Julia running in parallel, with the bootstrap data drawn groupwise from groups defined by year. One could attain the same results with bootstrap:reghdfe or bs:reghdfe. But reghdfejl implements vce(bs) in a radically faster way. After launching the extra instances of Julia to split the work, it copies the primary data just once to each of them, and effects the sampling in Julia by updating a weight variable (cluster 1’s weights are doubled if drawn two times in a bootstrap sample, and so on.).
To demonstrate reghdfejl’s speed, I create a data set with 10 million rows and two sets of fixed effects. I run the same regression with reghdfe and reghdfejl, as well as with the built-in areg command, which in StataNow 18.5 gained the ability to absorb multiple fixed effects. While reghdfe and reghdfejl can cluster standard errors on multiple variables (CGM_2011), areg cannot, so the example is restricted to one-way clustering:
@@
    areg      y x1 x2, a(id1 id2) cluster(id1)
    reghdfe   y x1 x2, a(id1 id2) cluster(id1) dof(none)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1) gpu
@@
    areg      y x1 x2, a(id1 id2) cluster(id1)
    reghdfe   y x1 x2, a(id1 id2) cluster(id1) dof(none)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1)
    reghdfejl y x1 x2, a(id1 id2) cluster(id1) gpu
For comparability, the reghdfe lines include dof(none) to disable the degrees-of-freedom correction, since reghdfejl lacks that feature.
The first quadruplet runs on one CPU core, as under a non-MP flavor of Stata. But that setting only affects the execution of reghdfe’s ado and Mata code: the Julia program still runs on multiple CPU threads. On a laptop with an Intel i9-13900H CPU and an NVIDIA RTX 4070 GPU, areg takes 8.1 seconds and reghdfe 17 seconds. reghdfejl takes 2.6 seconds without the gpu option and 1.8 with. Moving to 6 processors in Stata and Mata—since the Intel chip has 6 “performance” cores—lowers the times to 5.0, 10, 2.2, and 1.8 seconds.
@@
Using Stata’s bs prefix command with reghdfe takes 69 seconds. Switching to the parallel package of George Vega and Brian Quistorff in order to split the work among 12 copies of Stata [^8] See github.com/gvegayon/parallel. cuts the time to 19 seconds. reghdfejl takes just 1.9 seconds. But, as usual, the Julia function is slower the first time it is invoked in a Stata session: this one spins up 12 worker copies of Julia and triggers some just-in-time compilation in each. The workers remain active after completing their first task, so subsequent bootstrapping is fast as long as the number of workers reqested by the procs() suboption of the vce() option does not rise further.
The reghdfejl package includes a partialhdfejl command for absorbing (partialling) fixed effects out of other variables. partialhdfejl command lines must contain either a generate() option to name the variables that will hold the results, or a prefix() option for making the new variable names out of the old. If any of the new variables already exists, partialhdfejl will error, unless the replace option is also included. The following two runs produce the same point estimates:
@@ 3.2 Wild bootstrapping
Roodman et al. (2019) introduces the Stata package boottest for inference via the wild bootstrap. Among the methods boottest implements is the “wild restricted efficient” bootstrap (WRE), which is the Davidson and MacKinnon (2010) adaptation of the wild bootstrap for instrumental variables (IV) estimation. Since the publication of Roodman et al. (2019), boottest has been optimized in ways that especially benefit the WRE. Using the Frisch-Waugh-Lowell theorem, exogenous controls are now partialled out of the other variables using the same algebraic tricks that accelerate boottest’s wild bootstrap for ordinary least squares. The partialling-out accelerates the application of the user’s IV estimator despite its irreducible nonlinearity (Roodman et al., 2019, section 6.1). Also incorporated (more fully), is MacKinnon (2023)’s point that the data only enter wild bootstrap computations through indexed sets of matrix products such as $\{{\bf X}^{\prime}_{g}{\bf Y}_{g}:g=1,...,G\}$ where ${\bf X}_{g}$ and ${\bf Y}_{g}$ hold the rows of data matrices ${\bf X}$ and ${\bf Y}$ for cluster $g$. Reducing the data to these products early in the calculation complicates the algebra the code must implement, but speeds execution when clusters are few.
And the functionality in boottest, so revised, has been transplanted to the Julia package WildBootTests.jl. Adding the julia option to a boottest command line switches to the new back end. [^9] Jan Ditzen suggested this interface for calling the Julia implementation.
Section 8.4 of Roodman et al. (2019) approximately replicates two 2SLS regressions in the Levitt (1996) study of the short-term impact of decarceration on crime. [^10] Data and preparation code are available at davidroodman.com/david/Levitt.zip. Here is a version of that example that substitutes reghdfejl for ivregress and sets the stage for calling boottest to perform inference about changes in the size of the prisoner population affect crime rates:
@@
On a single CPU core, version 2.3.5 of boottest, the one published with Roodman et al. (2019), needs 210 seconds to perform the two boottest calls. The latest version finishes in 7.4 seconds, even without calling Julia. Appending julia to the boottest line above lowers the time to 4.1 seconds. (However, the first time boottest is called after installation, there is a significant delay for pre-compilation, just as when jl is first run, as described in section section 2.1.)
The R package fwildclusterboot (Fischer and Roodman, 2021) also calls WildBootTests.jl to perform the WRE, and optionally does so to perform the wild bootstrap after ordinary least squares. This may be the first instance of a Julia statistical package serving as a back end on multiple platforms.
@@ 3.3 Plotting
Many data visualization packages are available for Julia. Surveying them is beyond the scope of this article. To indicate their potential for Stata users, I generate two graphs with the plotting package Makie.jl.
Makie works with several plotting back ends, which render output in formats designed for different applications, such as interactive graphs in web pages and publication-quality image files. The Cairo back end is best for the latter. It is bundled with Makie in CairoMakie.jl:
  • jl AddPkg CairoMakie  # one-time installation
    jl: using CairoMakie  # use now
In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.111111blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command. margins is then called to compute the probability of high blood pressure according to the fitted model, across a grid of ages and weights. The probabilities are copied to Python and depicted in a three-dimensional surface plot. An adaptation of that example to Julia:
@@
This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window, That is, I drop the “;” and “///” from the ends of lines. The first Julia line imports the data as a matrix, converts it to a DataFrame, and names the columns. Then the surface plot is built and stored in the Julia variable f. The call to surface() refers to columns of the DataFrame with the convenient dot syntax, as in df.age; here, column names are not prefixed with a colon. The next command displays the figure in a new window, which might pop up behind the Stata window. That line takes advantage of Julia’s piping syntax for function calls, using the symbol “|>”. (display(f) would be equivalent.) Last, the image is saved to disk. It appears in Figure 1.
@@
dropmissing!(df, :bmi) is similar to drop if bmi==. in Stata. However, it also changes the type of df.bmi from Float32? to Float32, meaning that the column no longer accepts missing values. That is apparently necessary to make Makie work. The “!” in “dropmissing!” has no special meaning in the syntax of Julia: it is just a character in the function’s name. But by convention, functions that modify their arguments have names ending in “!”.
The for...enumerate() loop iterates over the racial groups. r takes race names as strings while i is a counter starting from 1. In the Stata data set, the race column is coded with 1, 2, and 3. Those codes are displayed through a value label as “White”, “Black”, and “Other.” Since jl save stores race as a CategoricalVector, the Julia column’s contents are outwardly strings, not numbers. df.bmi[df.race .== r] uses Boolean indexing to extract the elements of the bmi column for which race takes the string value in r.
Each iteration of the loop plots a density in slate gray with 40% opacity. The vertical spacing offset of $1/20$ is interpreted in the vertical axis’s units, BMI density, and was chosen by trial and error to cause overlapping. See Figure 2.
@@ 4.1 The JuliaMono font
Julia output routinely contains certain Unicode characters, such as vertical and diagonal ellipses, that Stata fonts do not cover. (However, this is not necessarily the case for Stata packages that call Julia back ends.) The JuliaMono font is a free, monospaced font that includes those characters, and looks reasonably good in Stata’s results window. Without it, some Julia output will appear slightly garbled. The font can be installed from github.com/cormullion/juliamono/releases. Type help fonts in Stata to learn how to change the results window font.
@@ 4.2 Multithreading
Julia includes several facilities for multitasking. One is distributed processing, which involves launching multiple copies of Julia that can pass messages and data to each other. reghdfejl’s vce(bs) option uses that feature. A lighter form is multithreading, in which multiple execution lines run within one Julia process, largely sharing the same data space. That is the sort @tturbo exploits. The maximum number of threads Julia allows is set when it launches, and it can default to 1. A good guess for the optimum is the number of “performance” or “big” cores on the CPU. You can determine the current limit with the command jl:Threads.nthreads().
How best to set Julia’s thread limit depends on the operating system. In Windows, you can open the Environment Variables control panel and assign JULIA_NUM_THREADS. In Linux, you can add an entry such as export JULIA_NUM_THREADS = 6 to $\sim$/.bashrc” or the equivalent. In macOS, environment variables do not affect graphical applications such as Stata when launched from the desktop. One can instead set the limit in each Stata session through the threads() option of the jl start command. Example: jl start, threads(8). This must be the first jl command issued after starting Stata. It must also precede invocation of Julia-dependent programs such as reghdfejl. [^12] In macOS, one can in fact assign JULIA_NUM_THREADS in $\sim$/.zshenv”. This will take effect when one launches Stata from the terminal. For example open -a StataMP will launch Stata/MP.
4.3 Limitations in accessing Stata objects
The julia package lets you transfer information between Stata and Julia with commands issued in either language; section 5 will lay out more details. However initiated, the job of copying falls to the C++ plugin included in the julia package. To read and write Stata macros, scalars, matrices, and variables, the plugin must in turn call the official Stata plugin interface. [^13] See stata.com/plugins. The SPI imposes certain constraints. For one, it makes public no functions that can create Stata variables or matrices. As a result, you can edit existing Stata variables and matrices from Julia, but not create new ones. The Stata-side jl subcommands can and do evade this constraint because they start from the Stata environment. For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy into it from a Julia matrix.
A subtler limitation pertains to local macros. Explaining it requires a deep dive. The “jl.ado” file, also part of the julia package, defines the jl command that is invoked repeatedly in the above examples. In order to call the C++plugin, it first declares the plugin’s existence with this line: